OpenStudy (anonymous):

Ok I am not sure if this was a mistake but I found it on a practice exam I was studying and it looks really hard to do: Integral of (25x(5x+1)^30) Expanding everything out would take way too long and I see no way to integrate that via u-substitution.

OpenStudy (anonymous):

i would try integrating by parts , with u=25x and dv=(5x+1)^30

OpenStudy (anonymous):

ok thanks, this is a calculus I course so idk why it was on the test then since i thought integration by parts was a calc II topic

OpenStudy (anonymous):

you havent covered integration by parts?

OpenStudy (anonymous):

I learned it back in high school, but integration by parts is not a topic that should be in calc I

OpenStudy (anonymous):

um then i dunno how else would u solve this other than integrating by parts

OpenStudy (anonymous):

the problem looks crazy if you input it in wolframalpha

OpenStudy (anonymous):

they have it expanded there thats why

OpenStudy (anonymous):

what i get for solution is \[\frac{ 5x }{ 31 }(5x+1)^{31}-\frac{ (5x+1)^{32} }{ 992 }\]

OpenStudy (amistre64):

\[25x(5x+1)^{30}\] \[25x\sum_{n=0}^{30}\binom{30}{n}(5x)^{30-n}(1)^{n}\] \[5^2x^1\sum_{n=0}^{30}\binom{30}{n}(5)^{30-n}~(x)^{30-n}\] \[\sum_{n=0}^{30}\binom{30}{n}(5)^{32-n}~(x)^{31-n}\] integrating part by part .... we get \[\sum_{n=0}^{30}\binom{30}{n}(5)^{32-n}~\frac{(x)^{31-n+1}}{31-n+1}\] seems to be something like that is my idea

OpenStudy (anonymous):

Figured it out. Set u=5x+1 and integrate it via u-substiturion. To solve the x probelm set x= u-1/5

OpenStudy (amistre64):

well yeah, if you want to be uncreative ... i guess thats a route you can take :)